OpenStudy (anonymous):

Solve the matrix equation AX = B for the 2  3 matrix X, given...

OpenStudy (anonymous):

\[A= \left[\begin{matrix}4 & -1 \\ -1 & 0\end{matrix}\right] and B=\left[\begin{matrix}3 & 7 & -2 \\ 0 & -4 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

I don`t know how to do this

OpenStudy (anonymous):

do I times each by the inverse of matrix A?

jimthompson5910 (jim_thompson5910):

yes you need to find the inverse of matrix A first then you multiply both sides on the left by this inverse

jimthompson5910 (jim_thompson5910):

so if C is the inverse of matrix A, then X = C*B

jimthompson5910 (jim_thompson5910):

keep in mind that the order matters and C*B is NOT the same as B*C

OpenStudy (anonymous):

so it should equal \[x= \left[\begin{matrix}0 & 4 & 1 \\ -3 & 9 & 6\end{matrix}\right]\]

OpenStudy (anonymous):

hey! matrix people! i need help with 1 more question! when you all get done here come on over to my problem. ;) http://openstudy.com/study#/updates/512d6933e4b098bb5fbc2976

jimthompson5910 (jim_thompson5910):

very nice KatClaire, you got it

OpenStudy (anonymous):

Thank you :)!!

jimthompson5910 (jim_thompson5910):

yw